# UniSage: A Unified and Post-Analysis-Aware Sampling for Microservices

### v1

@@ 4.2. Unified Data Fusion and Correlation
The foundational phase of the UniSage is to transform raw, disconnected telemetry streams into a unified and contextually rich representation. This phase is crucial as it establishes the instance-level correlations between traces and logs, enabling the sophisticated cross-modal analysis that follows. The process involves two main stages: initial pre-processing of each telemetry type and the subsequent correlation of traces and logs via their shared trace_idtrace\_id.
@@ 4.2.1. Telemetry Pre-processing
  • Log Parsing: The raw log streams, denoted as a set $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$, are typically parsed into a structured format for automated analysis (jiang2024large, ; jiang2024lilac, ). We employ Drain (drain, ), a widely-used online log parsing tool, to convert unstructured log messages into a set of parameterized templates $TP=\{tp_{1},tp_{2},\ldots,tp_{m}\}$. Each log entry $l_{i}$ is thus represented by its template, timestamp, service of origin, and verbosity level (e.g., INFO, WARN, ERROR).
  • Trace Encoding: Distributed traces, which record the end-to-end execution of requests, are inherently structured as directed acyclic graphs (DAGs). Each trace $t\in\mathcal{T}$ is composed of spans, where each span represents an operation within a service. To facilitate topological analysis, we decompose each trace graph $G_{t}$ into a set of unique root-to-leaf call paths, $\mathcal{C}_{t}=\{c_{1},c_{2},\ldots,c_{k}\}$. An individual call path $c_{j}$ is an ordered sequence of operations $O$, where each call path is an ordered sequence of operations $(s,o)$.
  • Metric Aggregating: Metrics are collected as multivariate time series data from each service. These metrics, denoted by $\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, are aggregated into fixed-time intervals (e.g., one minute) to form vectors that represent the service’s state at discrete points in time.
@@ 4.2.2. Cross-Modal Context Correlation
We assume a modern observability setup where logs generated during the lifecycle of a request are instrumented with the corresponding trace_idtrace\_id. This shared identifier is the key to our unified analysis. Before the main analysis pipeline, we perform a stream-processing step that establishes a direct link between each trace and its associated logs. Formally, for any given trace_idtrace\_id, denoted $t^{id}$, we define a correlation function $C$ that maps the identifier to a contextual tuple: $C(t^{id}\to(t,\mathcal{L}_{t})$, where $t$ is the unique trace corresponding to $t^{id}$, and $\mathcal{L}_{t}$ is the set of all log messages that share the same identifier. This creates a transient mapping for each active request, linking its complete distributed trace with the full set of logs generated during its execution across all services. This unified and correlated context is fundamental for the subsequent cross-modal signal propagation bias (Section 4.4.2).
@@ 4.3.1. Model-wise Anomaly Detection
The first stage of our analysis is a continuous process of anomaly detection performed independently for each data modality. These detectors are designed to be lightweight, serving as an efficient first pass to identify preliminary signals of system malfunction without incurring significant computational overhead.
  • Log-wise Detection: Building on the parsed log templates $tp_{i}\in TP$ from the pre-processing phase, the log-wise detector focuses on significant shifts in log event frequencies. We primarily monitor the rates of ERROR and WARN level logs per service. We apply the simple yet effective $k$-$\sigma$ rule to detect anomalies. For each service, we compute a weighted log rate combining error and warning logs: $r=w\cdot r^{\text{e}}+(1-w)\cdot r^{\text{w}},$ where $w\in[0,1]$ is a tunable parameter. We set $w$ = 0.8 since error logs are generally more indicative of failures. An anomaly is flagged if this rate deviates from the fault-free mean $\overline{r}$ by more than $k$ standard deviations: $r-\overline{r}>k\cdot\sigma.$
  • Trace-wise Detection: For distributed traces, we use span latency as the primary indicator of performance anomalies. Similar to the log-wise detector, we apply the k-sigma rule to the duration of spans within each service. Services associated with traces containing spans that exhibit abnormal latency are marked as potentially anomalous for further investigation.
@@ 4.3.2. Root Cause Analysis
Once an anomaly is detected, the root cause analysis (RCA) phase further investigates the anomaly and produces a ranked list of root causes, along with interpretable diagnostic reports for SREs to mitigate it. Unlike conventional approaches that sample data before fault analysis, potentially risking the loss of critical information, UniSage conducts RCA on the complete stream of observability data. This design helps balance diagnostic accuracy with storage efficiency. The analysis consists of two steps. First, the Pattern mining identifies deviations from normal behavior and generates pattern events. Each event includes both normal and abnormal values, along with the specific pattern of deviation, allowing SREs to quickly pinpoint anomalous behavior. Then, the Anomaly service scoring prioritizes suspected root cause services based on the observed anomalies.
Table 1. Multi-modal Pattern Templates and Examples.
Modality Pattern Template Example
Log $(ranking,logTemp,service,normal$ $Freq,observedFreq,deviation)$ $(1,GET/api/v1/consignservice/$ $consigns/\dots 502,frontend,0,4.05,4.05)$
@@
Metric $(ranking,metricName,service,normal$ $Avg,observedAvg,deviation)$ $(1,k8s.pod.cpu.usage,$ $consign,0.015,0.241,381.333)$
Trace $(ranking,spanName,service,normal$ $Lat,observedLat,deviation)$ $(1,Session.mergeconsign.entity$ .consignRecord,consign,0.03,0.56,17.91).consignRecord,consign,0.03,0.56,17.91)
Pattern mining.
@@
To handle the heterogeneity and high volume of multi-modal observability data, we design a lightweight pattern miner to extract informative patterns from logs, traces, and metrics. Instead of processing telemetry data streams directly, the miner transforms them into a unified, structured representation of patterns, as illustrated in Table 1. This design ensures processing efficiency and high recall, enabling real-time monitoring without introducing significant latency. For each modality, the miner quantifies the deviation of current behavior from fault-free reference data:
@@
  • Metric Pattern. We generate metric patterns by quantifying the deviation of their current behavior from a known fault-free reference period. We measure the change between the current distribution $Q$ and the reference distribution $P$ using two indicators: the distributional overlap $\mathcal{O}=\int\min(P(x),Q(x))\,dx$, and the relative mean value shift $\Delta^{\mu}$. These indicators are combined into a composite deviation score $\Delta^{\mathcal{M}}=k\cdot\mathcal{O}+(1-k)\cdot\Delta^{\mu},$ where the weighting factor $k$ is fixed at 0.8. This score provides a unified measure that captures both the stability of the metric distribution and the shift in its central tendency.
@@
  • Trace Pattern. For distributed traces, patterns identify end-to-end requests with significant performance degradation. A deviation score $\Delta^{\mathcal{T}}$ is calculated for each trace, representing the maximum relative latency increase of any of its constituent spans compared to its reference latency: $\Delta^{\mathcal{T}}(t)=\max_{sp_{i}\in t}(\frac{lat(sp_{i})-\widehat{lat}(sp_{i})}{\widehat{lat}(sp_{i})})$.
@@
  • Log Pattern. Log patterns are derived from significant increases in ERROR message frequencies. For each ERROR log template $l$, the deviation score $\Delta^{\mathcal{L}}(l)$ is defined as $\Delta^{\mathcal{L}}(l)=\frac{r(l)-\hat{r}(l)}{\hat{r}(l)+\epsilon},$ where $r(l)$ is the current frequency, $\hat{r}(l)$ is the fault-free reference frequency, and $\epsilon$ is a small constant to ensure stability for unseen templates (where $\hat{r}(l)=0$).
Anomaly service scoring.
@@
Based on the mined patterns, we compute a real-time anomaly score for each service $s$, which reflects its likelihood of being the root cause of an issue. Let $A_{s}^{\mathcal{Z}}$ be the set of patterns of modality $\mathcal{Z}$ associated with service $s$. Each pattern $a\in A_{s}^{\mathcal{Z}}$ has a deviation score $\Delta_{a}^{\mathcal{Z}}$ and a rank $R_{a}^{\mathcal{Z}}$ (used to prioritize more severe deviations). The service-level anomaly score is defined as:
@@
where $|A_{s}^{\mathcal{Z}}|$ is the number of patterns of modality $\mathcal{Z}$ for service $s$, and $|A^{\mathcal{Z}}|$ is the total number of patterns of that modality across all services. This produces a ranked list of services, which serve as the input for anomalous service bias to guide the analysis-guided sampler (Section 4.4.2) in the subsequent phase.
@@ 4.4. Dual-Pillar Sampling Strategy
The final phase of the UniSage framework is to synthesize all available information into a robust and efficient sampling decision. We introduce a dual-pillar strategy that decomposes the sampling logic into two complementary components: a Edge-case-based Sampler that provides a reference of coverage for rare and unusual data points, and an Analysis-guided Sampler that precisely targets data related to detected anomalies. This structure ensures that UniSage is both sensitive to known failure patterns and resilient to novel ones.
@@ 4.4.1. Edge-case-based Sampler
Topological Bias.
This bias captures structural deviations in distributed traces. A reference set of call paths $C_{\mathcal{F}}$ is first constructed from fault-free traces. For an incoming trace $t$, we extract its call paths $C_{\mathcal{P}}$ and assess their dissimilarity to the reference set. Specifically, the similarity between a path $c\in C_{\mathcal{P}}$ and a reference path $c^{\prime}\in C_{\mathcal{F}}$ is measured via the Jaccard similarity: $\text{Sim}(c,c^{\prime})=\frac{|c\cap c^{\prime}|}{|c\cup c^{\prime}|}.$ A score of 1 denotes an identical match, while 0 indicates a complete structural deviation. The topological bias score for trace $t$ is then defined as: $Score_{topo}(t)=1-\max_{c^{\prime}\in C_{\mathcal{F}}}\text{Sim}(c,c^{\prime}).$ To reduce computation, if a path $c$ is identical to a reference path, we directly assign $\text{Sim}(c,c^{\prime})=1$ and omit further comparisons. This pruning strategy substantially improves efficiency while preserving accuracy. For interpretability, when $Score_{topo}(t)>0$, the path $c$ and its closest reference $c^{\prime}$ are reported as evidence of structural novelty.
Behavioral Bias.
This bias evaluates the statistical rarity of observed behaviors, including span latencies in traces and template frequencies in logs. For Traces, this bias captures performance anomalies reflected in span latencies. For each span $sp_{i}$ within a trace $t$, we compute its standardized deviation (z-score) with respect to the fault-free distribution of span durations: $z(sp_{i})=\frac{\left|lat(sp_{i})-\mu\right|}{\sigma},$ where $\mu$ and $\sigma$ denote the mean and standard deviation under normal conditions. The behavioral bias score of a trace is defined as the maximum span-level deviation: $Score_{b}(t)=\max_{sp_{i}\in t}z(sp_{i}).$ A higher value indicates that the trace contains at least one span with significantly abnormal latency. Finally, the edge-case sampling probability is determined by combining the behavioral and topological bias scores:
@@
For Logs, this bias quantifies the rarity of log templates. Let $p_{l}^{\prime}$ denote the empirical probability of template $l$ observed during the fault-free phase $\mathcal{F}$. The behavioral bias score of a log is defined as: $Score_{b}(l)=-\log(p_{l}^{\prime}+\epsilon),$ where $\epsilon$ is a small constant for numerical stability (e.g., $10^{-6}$). Templates unseen in $\mathcal{F}$ are treated as maximally rare by setting $p_{l}^{\prime}=0$. The corresponding sampling probability is then normalized using the same scaled sigmoid function: $p_{e}(l)=\frac{2}{1+e^{-2Score_{b}(l)}}-1.$ This approach ensures that both unseen and low-frequency templates are prioritized during sampling.
@@ 4.4.2. Analysis-guided Sampler
Anomalous Service Bias.
This bias incorporates the final output of the root cause analysis (Section 4.3.2). All data instances (traces and logs) originating from or passing through services with a high anomaly score $\varphi_{s}$ are assigned a proportionally higher sampling probability $p_{a}$. This ensures that data surrounding the most likely source of a failure is comprehensively retained.
For Traces. Given a trace $t$, its analysis-guided sampling probability $p_{a}(t)$ is determined by the maximum anomaly score among all services on its execution path. This is because the most significant anomaly often originates from a specific root cause service. We use the final service score $\varphi_{s}$ from the RCA stage. The probability is defined as: $p_{a}(t)=\frac{2}{1+e^{-2\max_{s\in\mathcal{S}_{t}}\varphi_{s}}}-1,$ where $\mathcal{S}_{t}$ is the set of services involved in trace $t$.
For Logs. Given a log $l$ associated with service $s$, its sampling probability is directly determined by the anomaly score of that service: $p_{a}(l)=\frac{2}{1+e^{-2\varphi_{s}}}-1.$
Cross-Modal Signal Propagation Bias.
This bias leverages the trace–log correlation established in Section 4.2.2 to propagate anomaly evidence across modalities. This ensures that evidence from one signal type can elevate the perceived importance of a correlated signal in another. We update the sampling probabilities $p_{a}(l)$ and $p_{a}(t)$ to reinforce cross-modal diagnostic signals.

### v2

@@ 3.2. Unified Data Fusion and Correlation
The first phase of the UniSage is to transform raw, disconnected telemetry streams into a unified and contextually rich representation. This phase is crucial as it establishes the instance-level correlations between traces and logs, enabling the sophisticated cross-modal analysis that follows. The process involves two main stages: initial pre-processing of each telemetry type and the subsequent correlation of traces and logs via their shared traceID.
@@ 3.2.1. Telemetry Pre-processing
Trace Encoding. Raw traces are not directly machine-readable and must be encoded into a machine-friendly representation to support downstream analysis such as anomaly detection, root cause analysis, and sampling (Huang et al., 2021, 2024a). As shown in Figure 1 (right), traces contain two complementary types of information that are particularly valuable: execution latency and topological structure. Therefore, our sampling strategy aims to preserve traces that exhibit abnormal execution durations or structural changes in their invocation topology. Based on this insight, we design a trace encoder consisting of two components: a duration-related encoder and a topology-related encoder. This dual representation enables both effective RCA and trace-aware adaptive sampling. The duration-related trace encoder parses tree-structured JSON traces into span-level records, where each span is encoded as a structured vector containing the traceID, spanID, span name, service, start time, and execution duration. These span-level statistics provide the primary signals for duration-based anomaly detection and RCA. The topology-related trace encoder captures structural information by decomposing each trace $t\in\mathcal{T}$ into a set of unique root-to-leaf call paths, denoted as $\mathcal{C}_{t}$. Each call path represents the service invocation order within the trace and is used to compute topological bias in subsequent analysis. To reduce redundancy, only unique call paths within each trace are retained.
Log Parsing. The raw logs $\mathcal{L}$ are typically parsed into a structured format for automated analysis (Jiang et al., 2024b, a). We employ Drain (He et al., 2017), a widely-used online log parsing tool, to convert unstructured log messages into a set of parameterized templates $TP=\{tp_{1},tp_{2},\cdots,tp_{k}\}$. Each log entry $l_{i}\in\mathcal{L}$ is thus represented by its template, timestamp, service, and verbosity level (e.g., INFO, WARN, ERROR).
Metric Aggregating. Metrics are collected as multivariate time series data from each service. These metrics are aggregated into fixed-time intervals (i.e., one minute) to form vectors that represent the service’s state at discrete points in time.
@@ 3.2.2. Cross-Modal Context Correlation
We assume a standard observability setup (i.e., OpenTelemetry (Foundation, 2025)) in which log messages generated during request execution are annotated with the corresponding traceID. Leveraging this shared identifier, we correlate traces with their associated logs before downstream analysis. Formally, for each trace $t$, we define a correlation function $g$ that maps it to a contextual tuple $g:t\rightarrow(t,\mathcal{L}_{t}),$ where $\mathcal{L}_{t}$ denotes the set of log messages associated with trace $t$. This correlation establishes a unified cross-modal context for each request, which serves as the foundation for cross-modal signal propagation bias (Section 3.4.2). When trace–log correlation is unavailable, UniSage disables cross-modal propagation while keeping the remaining pipeline intact.
@@ 3.3.1. Model-wise Anomaly Detection
The first stage of our analysis performs anomaly detection independently for each data modality. This modality-wise design is motivated by the observation that different fault types manifest distinctively across logs, traces, and metrics (Lee et al., 2023a, b). All detectors are intentionally lightweight, serving as an efficient first-pass filter that surfaces early abnormal signals with minimal computational overhead. Recent studies show that simple statistical methods often outperform more complex models in streaming and production settings (Fang et al., 2025). Therefore, for logs and traces, which are encoded as one-dimensional feature vectors, we adopt the $k-\sigma$ rule for anomaly detection. For metrics, which form high-dimensional time-series vectors, we apply PCA-based dimensionality reduction (Bourlard and Kamp, 1988) and detect anomalies using reconstruction error.
  • Log-wise Detection: Based on the parsed log templates obtained during preprocessing, the log-wise detector monitors abnormal shifts in log event frequencies at the service level. We focus on ERROR and WARN logs, which are more indicative of failures. For each service, we compute a weighted log rate $r=w\cdot r^{\text{e}}+(1-w)\cdot r^{\text{w}},$ where $r^{\text{e}}$ and $r^{\text{w}}$ denote the rates of error and warning logs, respectively, and $w\in[0,1]$ is a tunable parameter. We set $w=0.8$ to emphasize error logs. An anomaly is flagged if the log rate deviates from the fault-free mean $\overline{r}$ by more than $k$ standard deviations: $r-\overline{r}>k\cdot\sigma.$
  • Trace-wise Detection: For distributed traces, we treat span execution duration as the primary indicator of performance anomalies. Similar to log-wise detection, we apply the $k-\sigma$ rule to span durations within each service. An anomaly is flagged if span duration deviates from the fault-free mean by more than $k$ standard deviations.
@@ 3.3.2. Root Cause Analysis
Once an anomaly is detected, the root cause analysis (RCA) phase further investigates the anomaly and produces a ranked list of root causes, along with interpretable diagnostic reports for SREs to mitigate it. Unlike conventional approaches that sample data before fault analysis, potentially risking the loss of critical information, UniSage conducts RCA on the complete stream of observability data. This design helps balance diagnostic accuracy with storage efficiency. The analysis consists of two steps. First, the Feature profiler identifies deviations from normal behavior and generates feature events. Each event includes both normal and abnormal values, along with the specific deviation feature, allowing SREs to quickly pinpoint anomalous behavior. Then, the Anomaly service scoring prioritizes suspected root cause services based on the observed anomalies.
Table 1. Structure and Examples of Multi-modal Anomaly Features.
Modality Feature Template Example
[…]
@@
Metric rank, metric, service, normal_avg, observed_avg, deviation (1, k8s.pod.cpu.usage, consign, 0.015, 0.241, 381.333)
[…]
@@
Feature Profiling. To handle the heterogeneity and high volume of multi-modal observability data, we design a lightweight feature miner to extract informative features from logs, traces, and metrics. Instead of processing telemetry data streams directly, the miner transforms them into a unified, structured representation of deviation tuples, as illustrated in Table 1. This design ensures processing efficiency and high recall, enabling real-time monitoring without introducing significant latency. For each modality, the miner quantifies the deviation of current behavior from fault-free reference data:
@@
  • Metric Feature. We generate metric features by quantifying the deviation of their current behavior from a known fault-free reference period. We measure the change between the current distribution $Q$ and the reference distribution $P$ using two indicators: the distributional overlap $\mathcal{O}=\int\min(P(x),Q(x))\,dx$, and the relative mean value shift $\Delta^{\mu}$. These indicators are combined into a composite deviation score $\Delta^{\mathcal{M}}=k\cdot\mathcal{O}+(1-k)\cdot\Delta^{\mu},$ where the weighting factor $k$ is fixed at 0.8. This score provides a unified measure that captures both the stability of the metric distribution and the shift in its central tendency.
@@
  • Trace Feature. For distributed traces, features identify end-to-end requests with significant performance degradation. A deviation score $\Delta^{\mathcal{T}}$ is calculated for each trace, representing the maximum relative latency increase of any of its constituent spans compared to its reference latency: $\Delta^{\mathcal{T}}(t)=\max_{sp_{i}\in t}(\frac{lat(sp_{i})-\widehat{lat}(sp_{i})}{\widehat{lat}(sp_{i})})$.
@@
  • Log Feature. Log features are derived from significant increases in ERROR message frequencies. For each ERROR log template $l$, the deviation score $\Delta^{\mathcal{L}}(l)$ is defined as $\Delta^{\mathcal{L}}(l)=\frac{r(l)-\hat{r}(l)}{\hat{r}(l)+\epsilon},$ where $r(l)$ is the current frequency, $\hat{r}(l)$ is the fault-free reference frequency, and $\epsilon$ is a small constant to ensure stability for unseen templates (where $\hat{r}(l)=0$).
@@
Anomaly Service Scoring. Based on the mined features, we compute a real-time anomaly score for each service $s$, which reflects its likelihood of being the root cause of an issue. Let $A_{s}^{\mathcal{Z}}$ be the set of features of modality $\mathcal{Z}$ associated with service $s$. Each feature $a\in A_{s}^{\mathcal{Z}}$ has a deviation score $\Delta_{a}^{\mathcal{Z}}$ and a rank $R_{a}^{\mathcal{Z}}$ (used to prioritize more severe deviations). The service-level anomaly score is defined as:
@@
where $|A_{s}^{\mathcal{Z}}|$ is the number of features of modality $\mathcal{Z}$ for service $s$, and $|A^{\mathcal{Z}}|$ is the total number of features of that modality across all services. This produces a ranked list of services, which serve as the input for anomalous service bias to guide the analysis-guided sampler (Section 3.4.2) in the subsequent phase.
@@ 3.4. Dual-Pillar Sampling Strategy
The final phase of the UniSage framework is to synthesize all available information into a robust and efficient sampling decision. As illustrated in Figure 5, we introduce a dual-pillar strategy that decomposes the sampling logic into two complementary components: a Edge-case-based Sampler that provides a reference of coverage for rare and unusual data points, and an Analysis-guided Sampler that precisely targets data related to detected anomalies. This structure ensures that UniSage is both sensitive to known failure features and resilient to novel ones.
@@ 3.4.1. Edge-case-based Sampler
Topological Bias. This bias captures structural deviations in distributed traces. A reference set of call paths $C_{\mathcal{C}}$ is first constructed from fault-free traces. For an incoming trace $t$, we extract its call paths $C_{\mathcal{O}}$ and assess their dissimilarity to the reference set. Specifically, the similarity between a path $c\in C_{\mathcal{O}}$ and a reference path $c^{\prime}\in C_{\mathcal{C}}$ is measured via the Jaccard similarity: $\text{Sim}(c,c^{\prime})=\frac{|c\cap c^{\prime}|}{|c\cup c^{\prime}|}.$ A score of 1 denotes an identical match, while 0 indicates a complete structural deviation. The topological bias score for trace $t$ is then defined as: $Score_{topo}(t)=1-\max_{c^{\prime}\in C_{\mathcal{C}}}\text{Sim}(c,c^{\prime}).$ To reduce computation, if a path $c$ is identical to a reference path, we directly assign $\text{Sim}(c,c^{\prime})=1$ and omit further comparisons. This pruning strategy substantially improves efficiency while preserving accuracy. For interpretability, when $Score_{topo}(t)>0$, the path $c$ and its closest reference $c^{\prime}$ are reported as evidence of structural novelty.
Behavioral Bias. This bias evaluates the statistical rarity of observed behaviors, including span latencies in traces and template frequencies in logs. For Traces, this bias captures performance anomalies reflected in span latencies. For each span $sp_{i}$ within a trace $t$, we compute its standardized deviation (z-score) with respect to the fault-free distribution of span durations: $z(sp_{i})=\frac{\left|lat(sp_{i})-\mu\right|}{\sigma},$ where $\mu$ and $\sigma$ denote the mean and standard deviation under normal conditions. The behavioral bias score of a trace is defined as the maximum span-level deviation: $Score_{b}(t)=\max_{sp_{i}\in t}z(sp_{i}).$ A higher value indicates that the trace contains at least one span with significantly abnormal latency. Finally, the edge-case sampling probability is determined by combining the behavioral and topological bias scores:
@@
For Logs, this bias quantifies the rarity of log templates. Let $p_{l}^{\prime}$ denote the empirical probability of template $l$ observed during the fault-free phase $\mathcal{F}$. The behavioral bias score of a log is defined as: $Score_{b}(l)=-\log(p_{l}^{\prime}+\epsilon),$ where $\epsilon$ is a small constant for numerical stability (e.g., $10^{-6}$). Templates unseen in $\mathcal{C}$ are treated as maximally rare by setting $p_{l}^{\prime}=0$. The corresponding sampling probability is then normalized using the same scaled sigmoid function: $p_{e}(l)=\frac{2}{1+e^{-2Score_{b}(l)}}-1.$ This approach ensures that both unseen and low-frequency templates are prioritized during sampling.
@@ 3.4.2. Analysis-guided Sampler
Anomalous Service Bias. This bias incorporates the final output of the root cause analysis (Section 3.3.2). All data instances (traces and logs) originating from or passing through services with a high anomaly score $\varphi_{s}$ are assigned a proportionally higher sampling probability $p_{a}$. This ensures that data surrounding the most likely source of a failure is comprehensively retained. For Traces. Given a trace $t$, its analysis-guided sampling probability $p_{a}(t)$ is determined by the maximum anomaly score among all services on its execution path. This is because the most significant anomaly often originates from a specific root cause service. We use the final service score $\varphi_{s}$ from the RCA stage. The probability is defined as: $p_{a}(t)=\frac{2}{1+e^{-2\max_{s\in\mathcal{S}_{t}}\varphi_{s}}}-1,$ where $\mathcal{S}_{t}$ is the set of services involved in trace $t$. For Logs. Given a log $l$ associated with service $s$, its sampling probability is directly determined by the anomaly score of that service: $p_{a}(l)=\frac{2}{1+e^{-2\varphi_{s}}}-1.$
Cross-Modal Signal Propagation Bias. This bias leverages the trace–log correlation established in Section 3.2.2 to propagate anomaly evidence across modalities. This ensures that evidence from one signal type can elevate the perceived importance of a correlated signal in another. We update the sampling probabilities $p_{a}(l)$ and $p_{a}(t)$ to reinforce cross-modal diagnostic signals.